# Multi-Modal Recommendation Unlearning for Legal, Licensing, and Modality Constraints

### v2

@@ A.2 Challenges in Multi-modal Recommendation Unlearning
Ill-effects of sharding: 1 Partitioning data breaks the graph structure. Partitioning based on items restricts the system’s ability to recommend items across different partitions to users, and partitioning based on users leads to insufficient training data and poor data heterogeneity (Koch and Soll 2023). This causes a significant drop ($\sim 30\%$) in performance (You et al. 2024). 2 This adds unlearning overhead of extra computational cost and time at the time of training . 3 When items that need to be forgotten come from multiple shards, the worst case efficiency falls to the level of retraining from scratch. 4 The model’s initial setup, partitioned based on users during training, poses a challenge to implementing item-related data unlearning, as it cannot be readily reconfigured to partition data based on items. 5 Item-related data unlearning concurrently with user-related data unlearning is not possible.
Burden of aggregation: 1 Additional steps to aggregate performance across shards introduce significant overhead costs, with the burden increasing proportionally to the number of shards (Ramezani et al. 2021). 2 This adds unlearning overhead of extra computational cost and time at the time of inference. 3 The partition-aggregation framework is prohibitive for the models which have already been trained.

### v3

@@ A.2 Challenges in Multi-modal Recommendation Unlearning
Ill-effects of sharding: ❶ Partitioning data breaks the graph structure. Partitioning based on items restricts the system’s ability to recommend items across different partitions to users, and partitioning based on users leads to insufficient training data and poor data heterogeneity (Koch and Soll 2023). This causes a significant drop ($\sim 30\%$) in performance (You et al. 2024). ❷ This adds unlearning overhead of extra computational cost and time at the time of training . ❸ When items that need to be forgotten come from multiple shards, the worst case efficiency falls to the level of retraining from scratch. ❹ The model’s initial setup, partitioned based on users during training, poses a challenge to implementing item-related data unlearning, as it cannot be readily reconfigured to partition data based on items. ❺ Item-related data unlearning concurrently with user-related data unlearning is not possible.
Burden of aggregation: ❶ Additional steps to aggregate performance across shards introduce significant overhead costs, with the burden increasing proportionally to the number of shards (Ramezani et al. 2021). ❷ This adds unlearning overhead of extra computational cost and time at the time of inference. ❸ The partition-aggregation framework is prohibitive for the models which have already been trained.